# Multi-modal Bayesian Neural Network Surrogates with Conjugate Last-Layer Estimation

### v1

@@ 1 Introduction and Background
In this work, we design two multi-modal surrogates following the motivating example of Bayesian optimization (BO). BO is a gradient-free optimization method for expensive black-box objective functions [16]. The BO procedure begins with an initial design of input values to span the search space, from which a surrogate model is trained to approximate the objective function. Then, an acquisition function guides observation of the objective function at new points and the surrogate model is updated to include each additional observation. In this paper, we focus on the surrogate model for the objective function, which approximates the objective function at each step based on the expensive evaluations performed so far. Most commonly, the surrogate model is a Gaussian process that allows for statistically principled modeling and uncertainty quantification. However, other surrogate models have been explored, including neural networks [[, see]]li_study_2023 and random forests [[, e.g.,]]mccullough_high-throughput_2020, williams_enabling_2020, jayarathna_experimental_2024. Uncertainty quantification in the surrogate model is important for the BO task, as the acquisition function, which determines the sequential selection of new points to evaluate, seeks to balance exploration and exploitation [16, 15, 7]. Because surrogate models have broader applications than BO, we will refer to the modeled function as a quantity of interest and only use the term “objective function” when specifically referencing the motivating application of optimization.
In this paper, we introduce two novel multi-modal Bayesian neural network (BNN) surrogate models motivated by BO for multi-modal data. The remainder of this section contains necessary background for the methods used. Section 2 introduces the two proposed surrogate models. We leverage conditionally conjugate distributions in the network’s last layer in stochastic variational inference (SVI) and provide a method to perform this conjugate SVI estimation in the presence of partially missing observations (Section 3). We demonstrate improved prediction accuracy and uncertainty quantification compared to unimodal surrogate models for both scalar and time series data (Section 4). Finally, we summarize our results and propose future research directions (Section 5).
@@ 1.1 Multi-modal Learning
The category of multi-modal learning that is most relevant to our goal of a multi-modal surrogate model is multi-modal fusion. Returning to our motivating example, multi-modal surrogate models for BO are a multi-modal fusion problem in the sense that they incorporate data from multiple sources and multiple modalities to more accurately predict unobserved values of the quantity of interest.
@@ 1.3 Bayesian Neural Networks
Bayesian Neural Networks (BNNs) are neural networks where the parameters (weights and biases) are given prior distributions and are estimated via their posterior distributions conditional on observed data. In this paper, we focus on BNN surrogate models due to their use in multi-modal fusion problems, their flexibility, and their relationship to more conventional Gaussian process surrogate models in BO. [29] show that an infinitely-wide BNN with a single hidden layer and Gaussian priors is a Gaussian process with a neural covariance function. [21] later show that deep infinitely-wide BNNs are also Gaussian processes. However, by using finite BNNs instead of GPs or infinite BNNs, we avoid assuming stationarity of variance in the model or needing to calculate the neural covariance function. The use of a BNN also allows us to produce a posterior sample of the weights of the network, amortizing the cost of future evaluations of the surrogate model at unobserved locations.
@@
BNN posteriors can be estimated in two main ways: Markov chain Monte Carlo [28] and stochastic variational inference [[, SVI;]]blundell_weight_2015. MCMC estimates the posterior by producing auto-correlated samples that (as the number of samples approaches infinity) follow the desired posterior distribution. Despite being very commonly used for MCMC on BNN models, for large networks, the No U-Turn Sampler [[, NUTS,]]hoffman_no-u-turn_2014 struggles both in terms of speed and exploration of the posterior. Therefore SVI is often preferred.
@@
SVI approximates the posterior of the network parameters fundamentally differently to MCMC. While MCMC approximates the posterior distribution via a collection of samples from it, SVI approximates the posterior by finding the closest match from an analytical family of distributions. For a model with parameters, $\bm{\theta}$, and data, $\bm{x}$, we want to find a distribution $Q(\bm{\theta})\approx p(\bm{\theta}|\bm{x})$. The distribution, $Q$, is chosen to minimize the KL-divergence, $D_{KL}(Q\|p)$. In practice, the family for $Q$ is often chosen to be the so-called mean-field approximation: independent Gaussian distributions over each component of $\bm{\theta}$ (possibly after first transforming constrained parameters into unconstrained space). This choice makes the problem of optimizing $Q$ simple, and often this can lead to a very good approximation of $p(\bm{\theta}|\bm{x})$. However, for posterior distributions with strong dependence between parameters, the approximation can be poor [[, e.g.,]]wang_lack_2004. BNN posterior distributions do have strong dependence between parameters.
@@
Bayesian last-layer neural networks [[, BLL;]]lazaro-gredilla_marginalized_2010 are also commonly used [[, e.g.,]]snoek_scalable_2015, and can be thought of as a special variational approximation to a BNN posterior. For BLLs, weights and biases in all layers except the last layer are estimated with point estimates, while weights and biases in the last layer are estimated with a variational distribution. The variational distribution in the last layer is often comprised of independent Gaussian distributions; however, distributions with more complex dependence have been proposed [[, e.g.,]]li_multi-fidelity_2020, harrison_variational_2024. The estimation method we propose in Section 3 is, to the best of our knowledge, the first fully Bayesian variational BNN approximation proposed incorporating dependence through conjugate full conditional distributions in the last layer.
@@ 2 Proposed Multi-modal Surrogate Models
We propose two BNN-based multi-modal surrogate models: a joint model and a layered model. The models are influenced by different surrogate modeling and multi-modal learning methods, and are distinct in their use of data from different modalities. These models are trained on observations $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{n}$ of a main data modality, and auxiliary observations $\{(\bm{x}^{(m)}_{j},\bm{y}^{(m)}_{j})\}_{j=1}^{n_{m}}$ for $m=1,\dots,M$ from $M$ additional data modalities.
@@ 2.1 Joint Model
$\bm{\Sigma}$ $\sim p(\bm{\Sigma})$ (7)
$\bm{W}_{k}$ $\sim p(\bm{W}_{k}),\quad k=0,\dots,\ell$ (8)
$\bm{b}_{k}$ $\sim p(\bm{b}_{k}),\quad k=0,\dots,\ell$ (9)
where $\bm{y}^{\prime}=\begin{bmatrix}\bm{y}^{\top}&\bm{y}^{(1)\top}&\dots&\bm{y}^{(M)\top}\end{bmatrix}^{\top}$ is an output vector comprised of all modalities concatenated together, and $NN(\cdot;\cdot)$ is a BNN as defined in (2)-(4). We call this model the “joint model,” and show a simplified illustration of this model in Figure 1(a). Intuitively, we expect that this kind of model would learn representations of the input, $\bm{x}$, in its hidden layers that capture shared qualities of all data modalities simultaneously.
@@
Figure 1: Examples of the joint and layered model architectures for data with a scalar quantity of interest and one 3-dimensional vector auxiliary modality. In the joint model, the output, $\bm{\mu}^{\prime}$ models the mean for both the main response, $y$, and the auxiliary response, $\bm{y}^{(1)}$, as a single vector $\bm{\mu}^{\prime}=(\mu,\ \bm{\mu}^{(1)\top})^{\top}$. In the layered model, the mean of the auxiliary response, $\bm{\mu}^{(1)}$, is modeled by a separate BNN surrogate model and is then used as input to the main BNN surrogate model.
@@ 2.2 Layered Model
The layered model is an alternate approach that draws more from the structure of surrogate models for multi-fidelity BO, co-Kriging and hierarchical Kriging. In this model, the additional data modalities are used as predictors in a neural network surrogate model for the main modality. Because there may not be observations of the additional modalities at the same input values as the observations of the main modality, we use separate neural network surrogate models to predict unobserved values of the additional modalities with uncertainty. This results in the model,
@@
$\bm{\mu}^{(m)}$ $=NN(\bm{x};\bm{W}_{0}^{(m)},\dots,\bm{W}_{\ell}^{(m)},\bm{b}_{0}^{(m)},\dots,\bm{b}_{\ell}^{(m)}),\quad m=1,\dots,M$ (13)
$\bm{\Sigma}$ $\sim p(\bm{\Sigma}),\quad\bm{\Sigma}^{(m)}\sim p(\bm{\Sigma}^{(m)}),\quad m=1,\dots,M$ (14)
$\bm{W}_{k}$ $\sim p(\bm{W}_{k}),\quad\bm{W}_{k}^{(m)}\sim p(\bm{W}_{k}^{(m)}),\quad k=0,\dots,\ell\text{ and }m=1,\dots,M$ (15)
$\bm{b}_{k}$ $\sim p(\bm{b}_{k}),\quad\bm{b}_{k}^{(m)}\sim p(\bm{b}_{k}^{(m)}),\quad k=0,\dots,\ell\text{ and }m=1,\dots,M.$ (16)
We show a simplified example of this model in Figure 1(b).
@@
3 Variational Estimation of Our BNN Models
In this section, we describe a variational approximation to the posterior of BNN parameters that we use for the joint and layered multi-modal surrogate models. First, we consider the approximation for a single BNN, then describe the approximation in the presence of missing data. The multi-modal surrogate models and conditional last layer estimation are implemented using Pyro [2, 31], and are available in the Python package mmbo. [^1] Code will be made publicly available prior to publication.
@@ 3.1 Conjugate Last Layer Variational Estimation
Consider a BNN for regression, with $m$ input dimensions and $k$ output dimensions. Let all weights, biases, and other parameters prior to the last layer be called $\bm{\Phi}$, and the activations of the last layer of $h$ hidden nodes be $\bm{z}_{\ell}(\bm{x};\bm{\Phi})\in\mathbb{R}^{h}$ for input $\bm{x}\in\mathbb{R}^{m}$, as shown in in (3). Let the weights in the last layer be $\bm{W}\in\mathbb{R}^{h\times k}$, and the biases of the outputs be $\bm{b}\in\mathbb{R}^{k}$. Then, overall, the network can be expressed as
@@
has a known analytical form. We then use a conditional last layer variational approximation,
with $Q^{\prime}$ being independent Gaussian distributions. This distribution provides a better approximation to the true posterior, and simplifies the optimization problem by reducing the complexity of the distribution to be fit to just $Q^{\prime}$.
This form of variational posterior minimizes KL-divergence in the following sense. Consider two distributions, $P(\bm{\theta},\bm{\psi})=P(\bm{\theta})P(\bm{\psi}|\bm{\theta})$ and $Q(\bm{\theta},\bm{\psi})=Q(\bm{\theta})Q(\bm{\psi}|\bm{\theta})$, with densities $p$ and $q$, respectively. We consider $P$ fixed and $Q$ as an approximation to $P$, such as in variational inference. The KL divergence of $Q$ from $P$ is:
Term (B) in equation (27) is the expectation of $D_{KL}(Q(\bm{\psi}|\bm{\theta})\|P(\bm{\psi}|\bm{\theta}))$ with respect to $Q(\bm{\theta})$. Because $D_{KL}(Q(\bm{\psi}|\bm{\theta})\|P(\bm{\psi}|\bm{\theta}))\geq 0$ for all $\bm{\theta}$, term (B) is minimized when $D_{KL}(Q(\bm{\psi}|\bm{\theta})\|P(\bm{\psi}|\bm{\theta}))=0$ for all $\bm{\theta}$, which is only true when $Q(\bm{\psi}|\bm{\theta})=P(\bm{\psi}|\bm{\theta})$.
Therefore, when performing variational inference on a model with a subset of parameters, $\bm{\psi}$, whose full conditional distribution based on the other parameters, $\bm{\theta}$, and the data, $\mathcal{D}$, is analytically tractable, a better approximation to $P(\bm{\theta},\bm{\psi}|\mathcal{D})$ can always be achieved using an approximate distribution $Q(\bm{\theta})P(\bm{\psi}|\bm{\theta},\mathcal{D})$ instead of an arbitrary $Q(\bm{\theta},\bm{\psi})$, especially a separable $Q(\bm{\theta})Q(\bm{\psi})$.
For the BNN model, the conjugate priors are
@@
The conjugacy of the last layer of a BNN in regression problems was noticed before by [10], who use a Matrix-Normal-Inverse-Wishart distribution to estimate the parameters in the last layer of a Bayesian last layer neural network. However, they do not explore fully Bayesian estimation of the network with a conjugate last layer conditioned on the weight values in the previous layers as we have developed here.

### v2

@@ 1 Introduction and Background
In this work, we design two multi-modal surrogates following the motivating example of Bayesian optimization (BO). BO is a gradient-free optimization method for expensive black-box objective functions [8]. The BO procedure begins with an initial design of input values to span the search space, from which a surrogate model is trained to approximate the objective function. Then, an acquisition function guides observation of the objective function at new points and the surrogate model is updated to include each additional observation. In this paper, we focus on the surrogate model for the objective function, which approximates the objective function at each step based on the expensive evaluations performed so far. Most commonly, the surrogate model is a Gaussian process that allows for statistically principled modeling and uncertainty quantification. However, other surrogate models have been explored, including neural networks [[, see]]li_study_2023, foldager_role_2023, brunzema_bayesian_2024 and random forests [[, e.g.,]]mccullough_high-throughput_2020, williams_enabling_2020, jayarathna_experimental_2024. Uncertainty quantification in the surrogate model is important for the BO task, as the acquisition function, which determines the sequential selection of new points to evaluate, seeks to balance exploration and exploitation [8, 9, 3]. Because surrogate models have broader applications than BO, we will refer to the modeled function as a quantity of interest and only use the term “objective function” when specifically referencing the motivating application of optimization.
In this paper, we introduce two novel multi-modal Bayesian neural network (BNN) surrogate models motivated by BO for multi-modal data. The remainder of this section contains necessary background for the methods used. Section 2 introduces the two proposed surrogate models. We leverage conditionally conjugate distributions in the network’s last layer in stochastic variational inference (SVI) and provide a method to perform this conjugate SVI estimation in the presence of partially missing observations (Section 3). We demonstrate improved prediction accuracy and uncertainty quantification compared to unimodal surrogate models for both scalar and time series data (Section 4), and examine the performance of multi-modal models on real-world data (Section 5). Finally, we summarize our results and propose future research directions (Section 6).
@@ 1.1 Multi-modal Learning
The challenge within multi-modal learning that is most relevant to our goal of a multi-modal surrogate model is multi-modal fusion from the above list. Returning to our motivating example, multi-modal surrogate models for BO are a multi-modal fusion problem in the sense that they incorporate data from multiple sources and multiple modalities to more accurately predict unobserved values of the quantity of interest. In this work, we consider the special case of multiple data sources of the same nature (for instance all sources are time series), which is a common multi-modal setting in scientific data. We still refer to our modeling strategy as multi-modal.
@@ 1.3 Bayesian Neural Networks
Bayesian Neural Networks (BNNs) are neural networks where the parameters (weights and biases) are given prior distributions and the posterior distributions of the parameters are estimated on observed data. In this paper, we focus on BNN surrogate models due to their use in multi-modal fusion problems, their flexibility, and their relationship to more conventional Gaussian process surrogate models in BO. [19] show that an infinitely-wide BNN with a single hidden layer and Gaussian priors is a Gaussian process with a neural covariance function. [12] later show that deep infinitely-wide BNNs are also Gaussian processes. However, by using finite BNNs instead of GPs or infinite BNNs, we avoid assuming stationarity of variance in the model or needing to calculate the neural covariance function. The use of a BNN also allows us to produce a posterior sample of the weights of the network, amortizing the cost of future evaluations of the surrogate model at unobserved locations.
@@
BNN posteriors can be estimated either by an ensemble of samples, e.g., Markov chain Monte Carlo [18] and Deep Ensembles [11]; or by approximate closed-form distributions, e.g., MC-Dropout [4], Laplace approximations [15], Stochastic Weight Averaging Gaussian [16], and stochastic variational inference [[, SVI;]]blundell_weight_2015. MCMC estimates the posterior by producing auto-correlated samples that (as the number of samples approaches infinity) follow the desired posterior distribution. Despite being very commonly used for MCMC on BNN models, for large networks, the No U-Turn Sampler [[, NUTS,]]hoffman_no-u-turn_2014 struggles both in terms of speed and exploration of the posterior. Deep Ensembles [11] use multiple initialization points and optionally multiple bootstrap samples of data to produce a random ensemble of trained networks. With bootstrap samples of data, this ensemble can approximate the posterior distribution for uninformative priors.
MC-Dropout estimates uncertainty by using Dropout [25] at inference time in addition to training time. MC-Dropout thus approximates the posterior distributions of weights as scaled Bernoulli distributions. Laplace approximations use curvature at the loss function minimum to approximate the posterior as a Gaussian distribution. Stochastic Weight Averaging Gaussian also approximates the posterior as a Gaussian distribution, but uses points from the end of the standard training process using stochastic gradient descent instead of the loss function’s curvature to calculate the first and second moments.
@@
We focus on SVI for our task to take advantage of the network structure in specifying the approximate posterior distribution. SVI approximates the posterior of the network parameters by finding the closest match from an analytical family of distributions. For a model with parameters, $\bm{\theta}$, and data, $\bm{x}$, we want to find a closed-form distribution $Q(\bm{\theta})$, which minimizes the KL-divergence, $D_{KL}(Q\|p)$, from the true posterior distribution, $p(\bm{\theta}|\bm{x})$. Commonly, the family for $Q$ is chosen to be the so-called mean-field approximation: independent Gaussian distributions over each component of $\bm{\theta}$ (possibly after first transforming constrained parameters into unconstrained space). This choice makes the problem of optimizing $Q$ computationally tractable [31, 23], and has been applied to estimate posterior distributions in applications such as genomics [[, e.g.,]]carbonetto_scalable_2012 and language modeling [[, e.g.,]]blei_latent_2003. However, for posterior distributions with strong dependence between parameters, the approximation can be poor [[, e.g.,]]wang_lack_2004. BNN posterior distributions do have strong dependence between parameters.
@@
Bayesian last-layer neural networks [[, BLL;]]lazaro-gredilla_marginalized_2010 are also commonly used [[, e.g.,]]snoek_scalable_2015. For BLLs, weights and biases in all layers except the last layer are estimated with point estimates, while weights and biases in the last layer are treated as random variables. As a consequence, there is less variability in the estimation of a BLL than a fully Bayesian neural network, where all weights and biases are treated as random variables. The distribution of the weights and biases in the last layer is often comprised of independent Gaussian distributions; however, distributions with more complex dependence have been proposed [[, e.g.,]]li_multi-fidelity_2020, harrison_variational_2024. The estimation method we propose in Section 3 incorporates dependence through conjugate full conditional distributions in the last layer.
@@ 2 Proposed Multi-modal Surrogate Models
We propose two BNN-based multi-modal surrogate models: a joint model and a layered model. The models are influenced by different surrogate modeling and multi-modal learning methods, and are distinct in their use of data from different modalities. The joint model models values from all data sources as one combined vector output from a single BNN. The layered model is comprised of two BNNs, one modeling auxiliary sources and one modeling the main quantity of interest. Output from the network modeling the auxiliary sources, including predicted values between observations, is used as additional input to the main network. These models are trained on observations $\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{n}$ of a main data modality, and auxiliary observations $\{(\bm{x}^{(m)}_{j},\bm{y}^{(m)}_{j})\}_{j=1}^{n_{m}}$ for $m=1,\dots,M$ from $M$ additional data modalities.
@@ 2.1 Joint Model
${\bm{\Sigma}}^{-1}$ $\sim\mathrm{Wishart}(\nu_{0},\bm{V}_{0}),$ (7)
$\begin{bmatrix}\bm{b}_{\ell}&\bm{W}_{\ell}\end{bmatrix}|\bm{\Sigma}$ $\sim\mathrm{MatrixNormal}(\bm{0},\bm{\Lambda}_{0},\bm{\Sigma})$ (8)
$\bm{W}_{k}$ $\sim\mathrm{N}(0,1),\quad k=0,\dots,\ell-1$ (9)
$\bm{b}_{k}$ $\sim\mathrm{N}(0,1),\quad k=0,\dots,\ell-1$ (10)
where $\bm{y}^{\prime}=\begin{bmatrix}\bm{y}^{\top}&\bm{y}^{(1)\top}&\dots&\bm{y}^{(M)\top}\end{bmatrix}^{\top}$ is an output vector comprised of all modalities concatenated together, and $NN(\cdot;\cdot)$ is a BNN as defined in (2)-(4). We call this model the “joint model,” and show a simplified illustration of this model in Figure 1(a). The statistical parameters of this model are $\bm{\Sigma}$ and $\bm{b}_{k},\bm{W}_{k}$ for $k=0,\dots,\ell$. The model is estimated by estimating the posterior distributions of these parameters with closed form variational approximations. We expand on the specific choice of priors (7)-(10) and their significance to estimation in Section 3. Intuitively, we expect that this kind of model would learn representations of the input, $\bm{x}$, in its hidden layers that capture shared qualities of all data modalities simultaneously.
@@
Figure 1: Examples of the joint (left) and layered (right) model architectures for data with a scalar quantity of interest and one 3-dimensional vector auxiliary modality. In the joint model, the output, $\bm{\mu}^{\prime}$ models the mean for both the main response, $y$, and the auxiliary response, $\bm{y}^{(1)}$, as a single vector $\bm{\mu}^{\prime}=(\mu,\ \bm{\mu}^{(1)\top})^{\top}$. In the layered model, the mean of the auxiliary response, $\bm{\mu}^{(1)}$, is modeled by a separate BNN surrogate model and is then used as input to the main BNN surrogate model.
@@ 2.2 Layered Model
The layered model is an alternate approach that draws more from the structure of surrogate models for multi-fidelity BO, co-Kriging and hierarchical Kriging. In this model, the additional data modalities are used as predictors in a neural network surrogate model for the main modality. Because there may not be observations of the additional modalities at the same input values as the observations of the main modality, we use separate neural network surrogate models to predict unobserved values of the additional modalities with uncertainty. This results in a model comprised of $M$ auxilliary networks, for $m=1,\dots,M$:
$\bm{y}^{(m)}$ $\sim\mathrm{N}(\bm{\mu}^{(m)},\bm{\Sigma}^{(m)}),$ (11)
$\bm{\mu}^{(m)}$ $=NN(\bm{x};\bm{W}_{0}^{(m)},\dots,\bm{W}_{\ell}^{(m)},\bm{b}_{0}^{(m)},\dots,\bm{b}_{\ell}^{(m)}),$ (12)
${\bm{\Sigma}^{(m)}}^{-1}$ $\sim\mathrm{Wishart}(\nu_{0}^{(m)},\bm{V}_{0}^{(m)}),$ (13)
$\begin{bmatrix}\bm{b}_{\ell}^{(m)}&\bm{W}_{\ell}^{(m)}\end{bmatrix}|\bm{\Sigma}^{(m)}$ $\sim\mathrm{MatrixNormal}(\bm{0},\bm{\Lambda}_{0}^{(m)},\bm{\Sigma}^{(m)})$ (14)
$\bm{W}_{k}^{(m)}$ $\sim\mathrm{N}(0,1),\quad k=0,\dots,\ell-1$ (15)
$\bm{b}_{k}^{(m)}$ $\sim\mathrm{N}(0,1),\quad k=0,\dots,\ell-1,$ (16)
whose outputs feed into a main network,
@@
${\bm{\Sigma}}^{-1}$ $\sim\mathrm{Wishart}(\nu_{0},\bm{V}_{0}),$ (19)
$\begin{bmatrix}\bm{b}_{\ell}&\bm{W}_{\ell}\end{bmatrix}|\bm{\Sigma}$ $\sim\mathrm{MatrixNormal}(\bm{0},\bm{\Lambda}_{0},\bm{\Sigma})$ (20)
$\bm{W}_{k}$ $\sim\mathrm{N}(0,1),\quad k=0,\dots,\ell-1$ (21)
$\bm{b}_{k}$ $\sim\mathrm{N}(0,1),\quad k=0,\dots,\ell-1.$ (22)
We show a simplified example of this model in Figure 1(b). The statistical parameters of this model are $\bm{\Sigma}$, $\bm{b}_{k},\bm{W}_{k}$ for $k=0,\dots,\ell$, $\bm{\Sigma}^{(m)}$ for $m=1,\dots,M$, and $\bm{b}_{k}^{(m)},\bm{W}_{k}^{(m)}$ for $m=1,\dots,M$ and $k=0,\dots,\ell$. The model is estimated by estimating the posterior distributions of these parameters with closed form variational approximations. We expand on the specific choice of priors (13)-(16) and (19)-(22) and their significance to estimation in Section 3.
@@
3 Variational Estimation of Proposed BNN Models
In this section, we describe a variational approximation to the posterior of BNN parameters that we use for the joint and layered multi-modal surrogate models. We consider the approximation for a single BNN, then describe the approximation in the presence of missing data. Because each of our proposed models is composed of one or more of these BNNs, the models can be fit by simultaneously estimating all of their component BNNs. The multi-modal surrogate models and conditional last layer estimation are implemented using Pyro [2, 21], and are available in the Python package mmbo. [^2] Code will be made publicly available prior to publication.
@@ 3.1 Conjugate Last Layer Variational Estimation
Consider a BNN for regression as defined in (2)-(4), with $m$ input dimensions and $k$ output dimensions. Let all weights, biases, and other parameters prior to the last layer be called $\bm{\Phi}$, and the activations of the last layer of $h$ hidden nodes be $\bm{z}_{\ell}(\bm{x};\bm{\Phi})\in\mathbb{R}^{h}$ for input $\bm{x}\in\mathbb{R}^{m}$, as shown in in (3). Let the weights in the last layer be $\bm{W}\in\mathbb{R}^{h\times k}$, and the biases of the outputs be $\bm{b}\in\mathbb{R}^{k}$. Then, overall, the network can be expressed as
@@
has a known analytical form. For the BNN model, the conjugate priors are
@@
The conjugacy of the last layer of a BNN in regression problems was utilized by [6], who use a Matrix-Normal-Inverse-Wishart distribution to estimate the parameters in the last layer of a Bayesian last layer neural network (BLL, as defined previously). However, when using variational feature learning as we are, this previous work uses MAP estimates of $\bm{\Sigma}$ instead of the conjugate Inverse-Wishart posterior, and does not include a procedure for training using partially-missing response vectors, $\bm{y}$, as we describe in Section 3.2.
With the full conditional distributions in (32) and (33), we then use a conditional last layer variational approximation,
with $Q^{\prime}$ being independent Gaussian distributions. In other words, the posterior distribution of every weight, $w_{ijk}$, in the matrix $\bm{W}_{k}$ for $k=0,\dots,\ell-1$, and every bias, $b_{jk}$, in the vector $\bm{b}_{k}$ for $k=0,\dots,\ell-1$ is approximated as a normal distribution $\mathrm{N}(\mu_{w_{ijk}},\sigma_{w_{ijk}})$ or $\mathrm{N}(\mu_{b_{jk}},\sigma_{b_{jk}})$, respectively. The values $\mu_{w_{ijk}},\sigma_{w_{ijk}},\mu_{b_{jk}},\sigma_{b_{jk}}$ are found by minimizing the KL-divergence of the approximate distribution (38) to the true posterior, or equivalently, maximizing the ELBO as is the goal of variational inference. There are no values to be optimized in layer $k=\ell$, as the variational distribution of those weights, biases, and the covariance matrix $\Sigma$ are determined by (32) and (33) with parameters (34)-(37).
The conjugate form of variational posterior minimizes KL-divergence in the following sense. Consider two distributions, $P(\bm{\theta},\bm{\psi})=P(\bm{\theta})P(\bm{\psi}|\bm{\theta})$ and $Q(\bm{\theta},\bm{\psi})=Q(\bm{\theta})Q(\bm{\psi}|\bm{\theta})$, with densities $p$ and $q$, respectively. We consider $P$ fixed and $Q$ as an approximation to $P$, such as in variational inference. The KL divergence of $Q$ from $P$ is:
Term (B) in Equation (42) is the expectation of $D_{KL}(Q(\bm{\psi}|\bm{\theta})\|P(\bm{\psi}|\bm{\theta}))$ with respect to $Q(\bm{\theta})$. Because $D_{KL}(Q(\bm{\psi}|\bm{\theta})\|P(\bm{\psi}|\bm{\theta}))\geq 0$ for all $\bm{\theta}$, term (B) is minimized when $D_{KL}(Q(\bm{\psi}|\bm{\theta})\|P(\bm{\psi}|\bm{\theta}))=0$ for all $\bm{\theta}$, which is only true when $Q(\bm{\psi}|\bm{\theta})=P(\bm{\psi}|\bm{\theta})$. Therefore, when performing variational inference on a model with a subset of parameters, $\bm{\psi}$, whose full conditional distribution based on the other parameters, $\bm{\theta}$, and the data, $\mathcal{D}$, is analytically tractable, a better approximation to $P(\bm{\theta},\bm{\psi}|\mathcal{D})$ can always be achieved using an approximate distribution $Q(\bm{\theta})P(\bm{\psi}|\bm{\theta},\mathcal{D})$ instead of an arbitrary $Q(\bm{\theta},\bm{\psi})$, especially a separable $Q(\bm{\theta})Q(\bm{\psi})$.
Our BNN variational posteriors fit this framework with $\bm{\theta}:=\bm{\Phi}$, $\bm{\psi}=(\bm{W},\bm{b},\bm{\Sigma})$, and all distributions are conditioned on $\mathcal{D}$. Because the distribution (28) has a known closed form, using (38) as the variational distribution provides a better approximation to the true posterior than any other variational distribution using the same $Q^{\prime}(\bm{\Phi})$. It also simplifies the optimization problem by reducing the complexity of the distribution to be fit to just $Q^{\prime}$.